OpenStudy (liz_beauty45):

f you can read 40 pages in 2 hours, how many pages can you read in 5 hours? A) 20 B) 80 C) 100 D) 200

OpenStudy (anonymous):

c

OpenStudy (anonymous):

100 pages

OpenStudy (liz_beauty45):

thx

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Not only is giving direct answers not allowed on OpenStudy, but it also doesn't help the one who asked to learn something from the question. Solution: 1. KNOW WHAT IS GIVEN. \[\huge \frac{ 40~pages }{ 2~hours }= \frac{ x~ pages }{ 5~hours }\] 2. CROSS MULTIPLY AND SOLVE FOR X. \[\huge2x=(5)(40)\]\[\huge 2x=200\] \[\huge x=100\]

OpenStudy (liz_beauty45):

thx...